A diver of mass 65 kg takes off from the 10 m platform from a standing position by jumping upwards with a (near) vertical speed of 2.0 ms-1.

Compare and contrast the trajectories of a diver leaving the 10 m platform with

i)A running take-off with a horizontal velocity of 3.0 ms-1

ii) A running take-off with a velocity of 3.0 ms-1 with an angle of elevation of 60°


1
Expert's answer
2021-03-02T18:04:41-0500

1)


"x=3t,y=10+2t-0.5(9.8)t^2"


2)


"x=3\cos{60}t,y=10+3\sin{60}t-0.5(9.8)t^2"